OpenStudy (anonymous):

in the scalene triangle ABC measure of angle B = 45 and measure of angle C = 55. what is the order of the sides of in length from longest to shortest? A. AB, BC, AC B. BC, AC, AB c. AC, BC, AB D. BC, AB, AC

OpenStudy (ryan123345):

first you have to do the Pythagorean Theorem to find A do the Formula is A^2+B^2=C^2 you plug it in and get A^2+45^2=55^2 so then it equals A^2+2025=3025 subtract 2025 from each sid and you get A^2=1000 than you square root and get A=100

OpenStudy (ryan123345):

nvm im dumb just realized its angles :P

OpenStudy (ryan123345):